THE CM-, MM- AND SUBMM-WAVE SPECTRUM OF ALLYL ISOCYANIDE AND RADIOASTRONOMI-CAL OBSERVATIONS IN ORION KL AND THE PRIMOS LINE SURVEY

\underline{I. HAYKAL}, R. A. MOTIYENKO, L. MARGULÈS, and T. R. HUET, Laboratoire PhLAM, UMR8523 CNRS-Université Lille 1, F-59655 Villeneuve d’Ascq Cedex, France; P. ECIJA, E J. COCINERO, F BASTERRETXEA, J. A. FERNÁNDEZ, F. CASTANO, Departamento de Química Física, Facultad de Ciencia y Tenología, Universidad del País Vasco, Barrio Sarriena s/n, 48940 Leioa (Spain); B. TERCERO, J. CERNICHARO, Centro de Astrobiología (CSIC-INTA). Ctra de Ajalvir, Km 4, 28850 Torrejón de Ardoz, Madrid, Spain; A. LESARRI, Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, 47011 Valladolid (Spain); J. C. GUILLEMIN, Sciences Chimiques de Rennes -Ecole Nationale Supérieure de Chimie de Rennes -CNRS -35700 Rennes, France.

Last year we presented the first rotational analysis of the ground state of the two conformers of allyl isocyanide from 4 GHz to 905 GHz. The analysis of the rotational spectrum of the cis conformer of allyl isocyanide was extended. We resolved Coriolis interactions of a and b types between the excited vibrational states ν_{1} = 1 and ν_{2} = 1, calculated to be at 156 cm^{−1} (A′) and 167 cm^{−1} (A′′) respectively (MP2/aug-cc-pvtz), from 150 GHz to 600 GHz^{a}. Strong perturbations were observed in the 150-310 GHz range for low values of the quantum number K_{a} starting from K_{a} = 0, 1. The anharmonicities appeared as well at higher frequencies for larger quantum numbers. The two modes were fitted together with the SPFIT/SPCAT^{b} suite of programs and a set of Coriolis parameters was accurately determined. The fit contains more than 3000 lines up to J = 99 and K_{a} = 12 for both modes. We did not detect these species neither in the IRAM 30-m line survey of Orion KL nor in the PRIMOS survey towards SgrB2. Nevertheless, we provided upper limits to their column density in Orion KL. This work was supported by the CNES and the Action sur Projets de l’INSU, PCMI.

aI. Haykal et al. manuscript in preparation

bH. Pickett J. Mol. Spec.\underline{148}, 371-377, 1991.